Pedro and Joanna decide to try a problem.


A baseball is thrown from the top of a tall building with an initial velocity of 10.5 m/s from a height of h = 10 m above the ground.


Find its speed when it reaches the ground if its launch angle is 37°.

vf =___  m/s


Find its speed if it is launched horizontally.

vf = ___ m/s


1
Expert's answer
2021-09-28T11:14:12-0400

In the first case, we will use the following equations:


"h_\text{up}=\frac{v^2\sin^2\theta}{2g},\\\space\\
h_\text{down}=h_\text{up}+h=\frac{v^2\sin^2\theta}{2g}+h=12\text{ m}."

The vertical component of velocity:


"v_y=\sqrt{2gh_\text{down}}=15.4\text{ m/s},\\
v_x=v\cos\theta=8.39\text{ m/s}^2.\\\space\\
v=\sqrt{v_y^2+v_x^2}=17.5\text{ m/s}."

In the second case:

"v_y=\sqrt{2gh}=14\text{ m/s},\\
v_x=v\cos\theta=8.39\text{ m/s}^2.\\\space\\
v=\sqrt{v_y^2+v_x^2}=16.3\text{ m/s}."